OpenStudy (anonymous):

Jim and Ciji Shaw must pay $.82 per $100 of coverage on their home which is insured for $80,000. How much is their annual insurance premium?

OpenStudy (phi):

you could use a ratio: \[ \frac{0.82}{100} = \frac{x}{80000} \] to solve for x, multiply both sides by 80000

OpenStudy (anonymous):

So would you get 100x=65600?

OpenStudy (phi):

yes, if you cross multiply you get 100 x= 0.82 * 80000 100 x = 65600 to get x by itself, you have to divide both sides by 100 \[ \frac{100 x}{100} = \frac{65600}{100} \] x= 656 the annual insurance premium is $656

OpenStudy (phi):

notice you can get to the answer more quickly if you multiply both sides by 80000: \[ \frac{0.82}{100} = \frac{x}{80000} \\ \frac{0.82\cdot 80000}{100} = \frac{x\cdot 80000 }{80000} \\\frac{0.82\cdot 80000}{100} =x \] if you simplify the left side you will get 656.